# A note on one-variable theorems for NSOP

### v2

@@ 1 Introduction
Lachlan’s proof is short but rather convoluted, so in Section 2 we give what is hopefully a simpler proof.
@@
Let $T$ be the theory of the structure $(\mathbb{Q},B)$, where $B(x,y,z)$ means that either $x<y<z$ or $x>y>z$. Note that $\operatorname{tp}(a,b)\neq\operatorname{tp}(b,a)$ for any distinct singletons $a,b$, in any model of $T$. The formula $B(x;y,z)$ has the SOP. However, no $L$-formula $\varphi(x;y)$ with $|y|=1$ has the SOP. Otherwise, take an ascending chain
$\varphi(\mathbb{M};b_{0})\subsetneq\varphi(\mathbb{M};b_{1})\subsetneq\cdots$
Then $\operatorname{tp}(b_{0},b_{1})=\operatorname{tp}(b_{1},b_{0})$, so $\varphi(\mathbb{M};b_{1})\subsetneq\varphi(\mathbb{M};b_{0})$, which is absurd.
The circular order is another example, again because $\operatorname{tp}(a,b)=\operatorname{tp}(b,a)$ for any two distinct singletons $a,b$.
@@ 2 A simplified proof of Lachlan’s theorem
An $L(\mathbb{M})$-formula $\varphi(x,y)$ is NSOP iff the poset of $\varphi$-sets $\{\varphi(\mathbb{M},b):b\in\mathbb{M}^{y}\}$ has finite height.
If $P$ is a poset with finite height, let $\operatorname{ht}:P\to\mathbb{N}$ be the height function. Thus $\operatorname{ht}(a)\geq k$ iff there is a chain $a_{0}<a_{1}<\cdots<a_{k}=a$. If an $L(\mathbb{M})$-formula $\varphi(x;y)$ is NSOP and $b\in\mathbb{M}^{y}$, let $\operatorname{ht}_{\varphi}(b)$ denote the height of $\varphi(\mathbb{M};b)$ in the poset of $\varphi$-sets. Then
@@ 4.1 Motivation
We may as well take $n=2$. Then $\varphi$ determines a partial order $<$ on $M^{2}$ with infinite height. For any fixed $a\in M$, we can restrict this partial order to $\{a\}\times M$. The resulting partial order must have finite height, or some one-variable formula witnesses SOP. Then ($(\{a\}\times M,<)$ must be a finite union of antichains. We may as well require $\{a\}\times M$ to be an antichain. Similarly, we may as well require $M\times\{a\}$ to be an antichain.
@@ 4.3 Labeled switchboards
Unfortunately, the theory $T^{-}$ fails to have quantifier elimination, because the class of switchboards does not have the amalgamation property (by Remark 4.24 below). We must first work in an expanded language, construct its model companion, then relate it back to the original setting.
@@
We pronounce $x\uparrow\{y,z\}$ as “$x$ favors $\{y,z\}$” and $x\downarrow\{y,z\}$ as “$x$ disfavors $\{y,z\}$.” Officially, we regard labeled switchboards as structures in a language $L^{+}$ with a 4-ary relation symbol $<$ and two 3-ary relation symbols $\uparrow$ and $\downarrow$. Note that $L^{+}$ expands $L^{-}$.
@@ 4.4 Single-element free amalgamation
Let $M$ be a labeled switchboard. Let $S$ be a subset and let $a_{1},a_{2}$ be two distinct elements of $M$. Then $a_{1}$ and $a_{2}$ are freely amalgamated over $S$ if the following holds:
@@
    The new edge $\{a_{1},a_{2}\}$ is incomparable to every other element of $[S\cup\{a_{1},a_{2}\}]^{2}$.
@@
4.5 The model companion $T$
@@
Let $M_{0}$ be the Fraïsse limit of finite labeled switchboards. Let $T^{+}$ be the complete theory of $M_{0}$. By general machinery, the following holds:
@@ 4.7 The theory has the SOP
and making everything else in $[\mathbb{Z}]^{2}$ be incomparable. Embed $\mathbb{Z}$ into a model $(M,<)$ of $T^{-}$. Then the poset $([M]^{2},<)$ has an infinite chain, inherited from $([\mathbb{Z}]^{2},<)$. Therefore $T^{-}$ and the equivalent theory $T^{+}$ have the SOP.
@@ 4.8 One-variable formulas
Work in a monster model $\mathbb{M}$ of $T^{+}$.
@@
Let $B\subseteq\mathbb{M}$ be small, and let $a_{1},a_{2}\in\mathbb{M}^{1}$ be two elements with $a_{1}\equiv_{B}a_{2}$.
@@
If $a_{1},a_{2}$ are singletons and $a_{1}\equiv_{B}a_{2}$, then $\operatorname{tp}(a_{1},a_{2}/B)$ is symmetric if and only if $a_{1}a_{2}\equiv_{B}a_{2}a_{1}$. This can be seen by quantifier elimination—the only atomic formulas in $\operatorname{tp}(a_{1},a_{2}/B)$ beyond those in $\operatorname{tp}(a_{1}/B)\cup\operatorname{tp}(a_{2}/B)$ are the formulas
@@
Let $B$ be a finite subset of $\mathbb{M}$, let $p(x)$ be a complete 1-type over $B$ not realized in $B$, and let $q(x,y)$ be a complete 2-type over $B$ extending $p(x)\cup p(y)$.
@@
    Take $c_{0}c_{1}$ to be any realization of $p$. For $i\geq 2$, take $c_{i}$ such that $c_{i-1}c_{i}\models q$, then use Lemma 4.31 to move $c_{i}$ by an automorphism over $Bc_{i-1}$ to make $c_{i}$ and $c_{0}$ be freely amalgated over $Bc_{i-1}$.
@@
    Proceed by induction on $i$. When $i=1$, Case (b) holds vacuously. Suppose $i>1$. Then $c_{i}$ and $c_{0}$ are freely amalgamated over $Bc_{i-1}$. Because of the definition of free amalgamation (i.e., Conditions (i), (ii) of Definition 4.16), there must be $\{u,v\}\in[Bc_{i-1}]^{2}$ such that $\{c_{i},b_{i}\}>\{u,v\}>\{c_{0},b_{0}\}$. If $\{u,v\}\subseteq B$ then we are in Case (a). Otherwise, $c_{i-1}\in\{u,v\}$, so $\{u,v\}=\{c_{i-1},b_{i-1}\}$ for some $b_{i-1}\in B$. Then $\{c_{i-1},b_{i-1}\}>\{c_{0},b_{0}\}$, so by induction, one of the following holds:
@@
    Now suppose $i\geq|B|$. We must show that $\operatorname{tp}(c_{0},c_{i}/B)$ is distinguished. There are two points to check in Definition 4.27. We check the second point; the first is similar222…using a variant of the claim where we replace $>$ with $<$. The proof is identical, but we can’t just say it works “by symmetry” since the definition of free amalgamation broke the symmetry between up and down in the poset. Suppose $b_{0},b_{i}\in B$ are such that
@@
    By the pigeonhole principle, there are $j_{1}<j_{2}\leq i$ with $b_{j_{1}}=b_{j_{2}}=:b$. Then $\{c_{j_{1}},b\}>\{c_{j_{2}},b\}$, which contradicts the axioms of switchboards.
@@ Acknowledgments.
The author was supported by the Ministry of Education of China (Grant No. 22JJD110002). Nick Ramsey encouraged the author to write up this note.

### v4

@@ 1 Introduction
Here, $|x|$ denotes the length of the tuple of variables $x$. Lachlan’s proof is short but rather convoluted, so in Section 2 we give what is hopefully a simpler proof.
@@
Let $T$ be the theory of the structure $(\mathbb{Q},B)$, where $B(x,y,z)$ means that either $x<y<z$ or $x>y>z$. Note that $\operatorname{tp}(a,b)=\operatorname{tp}(b,a)$ for any distinct singletons $a,b$, in any model of $T$. The formula $B(x;y,z)$ has the SOP. However, no $L$-formula $\varphi(x;y)$ with $|y|=1$ has the SOP. Otherwise, take an ascending chain
$\varphi(M;b_{0})\subsetneq\varphi(M;b_{1})\subsetneq\cdots$
Then $\operatorname{tp}(b_{0},b_{1})=\operatorname{tp}(b_{1},b_{0})$, so $\varphi(M;b_{1})\subsetneq\varphi(M;b_{0})$, which is absurd.
The dense circular order is another example, again because $\operatorname{tp}(a,b)=\operatorname{tp}(b,a)$ for any two distinct singletons $a,b$.
@@ 1.1 Conventions
1.2 Outline
Section 2 gives a simplified proof of Lachlan’s theorem mentioned above. Section 3 builds off this to prove the one-variable theorem for NSOP in the NIP case. Section 4—the bulk of the paper—works through the theory of switchboards, whose model companion is an example of a theory as in Theorem 1.7.
  • Subsection 4.1 gives motivation for our choice of theory.
  • Subsection 4.2 introduces the theory of switchboards.
  • Subsection 4.3 introduces an expansion, the theory of labeled switchboards. Labeled switchboards will have the amalgamation property, unlike unlabeled switchboards.
  • Subsection 4.4 verifies most of the amalgamation property for labeled switchboards.
  • Subsection 4.5 builds off this to prove that the theory of labeled switchboards has a model completion $T^{+}$.
  • Subsection 4.6 shows that the original class of (unlabeled) switchboards has a model companion $T^{-}$, definitionally equivalent to $T^{+}$.
  • Subsection 4.7 checks that $T^{+}$ and $T^{-}$ have the SOP.
  • Finally, Subsection 4.8 verifies the property that $\varphi(x,y)$ is NSOP whenever $|y|=1$.
@@ 2 A simplified proof of Lachlan’s theorem
Recall that an $L(\mathbb{M})$-formula $\varphi(x,y)$ is NSOP iff the poset of $\varphi$-sets $\{\varphi(\mathbb{M},b):b\in\mathbb{M}^{y}\}$ has finite height. If $P$ is a poset with finite height, let $\operatorname{ht}:P\to\mathbb{N}$ be the height function. Thus $\operatorname{ht}(a)\geq k$ iff there is a chain $a_{0}<a_{1}<\cdots<a_{k}=a$. If an $L(\mathbb{M})$-formula $\varphi(x;y)$ is NSOP and $b\in\mathbb{M}^{y}$, let $\operatorname{ht}_{\varphi}(b)$ denote the height of $\varphi(\mathbb{M};b)$ in the poset of $\varphi$-sets. Then
@@ 4.1 Motivation
We may as well take $n=2$. Then $\varphi$ determines a partial order $<$ on $M^{2}$ with infinite height. For any fixed $a\in M$, we can restrict this partial order to $\{a\}\times M$. The resulting partial order must have finite height, or some one-variable formula witnesses SOP. Then ($(\{a\}\times M,<)$ must be a finite union of antichains. We may as well require $\{a\}\times M$ to be a single antichain. Similarly, we may as well require $M\times\{a\}$ to be a single antichain.
@@ 4.2 The definitions and goal
In practice, we will always regard $<$ as a binary relation on $[M]^{2}$, and never think of it as a binary relation on $M^{2}$.
@@ 4.3 Labeled switchboards
Unfortunately, the theory $T^{-}$ fails to have quantifier elimination, because the class of switchboards does not have the amalgamation property (by Remark 4.24 below). We must first work in an expanded language, construct its model companion, then relate it back to the original setting. In this section, we define the expanded language.
@@
We pronounce $x\uparrow\{y,z\}$ as “$x$ favors $\{y,z\}$” and $x\downarrow\{y,z\}$ as “$x$ disfavors $\{y,z\}$.” Officially, we regard labeled switchboards as structures in a language $L^{+}$ with a 4-ary relation symbol $<$ and two 3-ary relation symbols $\uparrow$ and $\downarrow$. Note that $L^{+}$ expands $L^{-}$. In practice, we will always regard $\uparrow$,$\downarrow$ , ↓ as binary relations between $M$ and $[M]^{2}$, and never as binary relations between $M$ and $M^{2}$.
@@ 4.4 Single-element free amalgamation
Triangle relations are strict partial orders—they are transitive and irreflexive. Our strategy for amalgamating labeled switchboards will essentially be to freely amalgamate the triangle relations as partial orders. For this reason, we first review how strict partial orders can be amalgamated.
@@
Let $M$ be a labeled switchboard. Let $S$ be a subset and let $a_{1},a_{2}$ be two distinct elements of $M\setminus S$. Then $a_{1}$ and $a_{2}$ are freely amalgamated over $S$ if the following holds:
@@
    The edge $\{a_{1},a_{2}\}$ is incomparable to every other element of $[S\cup\{a_{1},a_{2}\}]^{2}$.
@@
4.5 The model companion $T^{+}$
We can now construct the model companion of the class of labeled switchboards.
@@ 4.5 The model companion $T^{+}$
Let $M_{0}$ be the Fraïsse limit of finite labeled switchboards. Let $T^{+}$ be the complete theory of $M_{0}$. By general machinery, the following hold:
@@ 4.6 Comparison to unlabeled switchboards
Next, we show that $T^{+}$ is a definitional expansion of some theory $T^{-}$, which is the model companion of unlabeled switchboards.
@@ 4.7 The theory has the SOP
and making everything else in $[\mathbb{Z}]^{2}$ be incomparable. Embed $\mathbb{Z}$ into a model $(M,<)$ of $T^{-}$. Then the poset $([M]^{2},<)$ has an infinite chain, inherited from $([\mathbb{Z}]^{2},<)$. Therefore $T^{-}$ and the definitionally equivalent theory $T^{+}$ (Section 4.5) have the SOP.
@@ 4.8 One-variable formulas
It remains to verify that $T^{\pm}$ is NSOP for formulas $\varphi(x,y)$ with $|y|=1$. Broadly speaking, the strategy is as follows. If such a formula $\varphi$ existed, there would be a definable strict partial order $(D,<)$ with $D\subseteq M^{1}$, and an ascending indiscernible sequence $a_{0}<a_{1}<a_{2}<\cdots$ in $(D,<)$. We build a new sequence $c_{0},c_{1},c_{2},\ldots$ in such a way that $c_{i}c_{i+1}\equiv a_{0}a_{1}$ for each $i$, but the $c_{i}$ are otherwise “indepedent” in some sense related to our earlier free amalgamation. By carefully analyzing different levels of independence, we will show that $c_{0}c_{n}\equiv c_{n}c_{0}$ for sufficiently large $n$. But $c_{0}<c_{1}<\cdots<c_{n}$, so it is impossible for $\operatorname{tp}(c_{0},c_{n})$ to equal $\operatorname{tp}(c_{n},c_{0})$.
For the remainder of the paper, work in a monster model $\mathbb{M}$ of $T^{+}$. Definitions 4.25 and 4.27 below should be seen as variants of “independence”, in some very loose sense.
@@
Let $B\subseteq\mathbb{M}$ be small, and let $a_{1},a_{2}\in\mathbb{M}\setminus B$ be two singletons with $a_{1}\equiv_{B}a_{2}$.
@@
If $a_{1},a_{2}\in\mathbb{M}\setminus B$ are singletons and $a_{1}\equiv_{B}a_{2}$, then $\operatorname{tp}(a_{1},a_{2}/B)$ is symmetric if and only if $a_{1}a_{2}\equiv_{B}a_{2}a_{1}$. This can be seen by quantifier elimination—the only atomic formulas in $\operatorname{tp}(a_{1},a_{2}/B)$ beyond those in $\operatorname{tp}(a_{1}/B)\cup\operatorname{tp}(a_{2}/B)$ are the formulas
@@
Let $B$ be a finite subset of $\mathbb{M}$, let $p(x)$ be a complete 1-type over $B$ not realized in $B$, and let $q(x,y)$ be a complete 2-type over $B$ extending $p(x)\cup p(y)\cup\{x\neq y\}$.
@@
    Take $c_{0}c_{1}$ to be any realization of $q$. For $i\geq 2$, take $c_{i}$ such that $c_{i-1}c_{i}\models q$, then use Lemma 4.31 to move $c_{i}$ by an automorphism over $Bc_{i-1}$ to make $c_{i}$ and $c_{0}$ be freely amalgated over $Bc_{i-1}$.
@@
    Proceed by induction on $i$. When $i=1$, Case (b) holds trivially. Suppose $i>1$. Then $c_{i}$ and $c_{0}$ are freely amalgamated over $Bc_{i-1}$. Because of the definition of free amalgamation (i.e., Conditions (i), (ii) of Definition 4.16), there must be $\{u,v\}\in[Bc_{i-1}]^{2}$ such that $\{c_{i},b_{i}\}>\{u,v\}>\{c_{0},b_{0}\}$. If $\{u,v\}\subseteq B$ then we are in Case (a). Otherwise, $c_{i-1}\in\{u,v\}$, so $\{u,v\}=\{c_{i-1},b_{i-1}\}$ for some $b_{i-1}\in B$. Then $\{c_{i-1},b_{i-1}\}>\{c_{0},b_{0}\}$, so by induction, one of the following holds:
@@
    Now suppose $i\geq|B|$. We must show that $\operatorname{tp}(c_{0},c_{i}/B)$ is distinguished. There are two points to check in Definition 4.27. We check the second point; the first is similar333…using a variant of the claim where we replace $>$ with $<$. The proof is identical, but we can’t just say it works “by symmetry” since the definition of free amalgamation broke the symmetry between up and down in the poset. Suppose $b_{0},b_{i}\in B$ are such that
@@
    By the pigeonhole principle, there are $j_{1}<j_{2}\leq i$ with $b_{j_{1}}=b_{j_{2}}=:b$. Then $\{c_{j_{1}},b\}>\{c_{j_{2}},b\}$, which contradicts the Switchboard Axiom.
@@ Acknowledgments.
The author was supported by the Ministry of Education of China (Grant No. 22JJD110002). Nick Ramsey encouraged the author to write up this note. The author would also like to thank the anonymous referee, who provided many helpful comments.